OpenStudy (anonymous):

Use the Pythagorean theorem to find the hypotenuse if one leg is 34 and the other leg is 18. Answer 28.84 38.47 35.56 28.48

OpenStudy (anonymous):

Remember the Pythagorean Theorem represents a way to solve a right triangle by relating its sides to the hypotenuse. Now, since a^2 + b^2 = c^2 we can assume a=34 and b=18. Now if we plug the two in the formula and set c^2 = to that number, then our job is almost done. The last step is to take the square root of both sides to get c= some number.

OpenStudy (anonymous):

Can SomeOne Help Me After This One :D